OpenStudy (anonymous):

okay so i already answered this question A. what is the density of an 84.7g sample of an unknown substance if the sample occupies 49.6cm^3? but question B. stumped me so question b. is what volume would be occupied by 7.75g of this same substance?

OpenStudy (anonymous):

well we know density=m/v

OpenStudy (anonymous):

and you solved a for m/v= and some density d

OpenStudy (anonymous):

and so now you have d=7.75/v so you are just solving for v

OpenStudy (anonymous):

because a substance's density stays the same even though the mass or the volume changes

OpenStudy (anonymous):

im still kinda confused about the d=7.75/v

OpenStudy (anonymous):

so you found the density in part a right?

OpenStudy (anonymous):

yeah

OpenStudy (anonymous):

ok what was that number

OpenStudy (anonymous):

1.71 g/cm^3

OpenStudy (anonymous):

ok so 1.71=7.75/v and you solve for v because a substance's density is constant and you are given the new mass so you just need to find v

OpenStudy (anonymous):

does that make sense?

OpenStudy (anonymous):

kinda so would the answer be like 4.53 g/cm^3

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

thank u soooo much for helping me

OpenStudy (anonymous):

np